What is the pOH of 0.022M NaC3H5O2?
*Please show every single step or I'll thumbs down. I really need to know how and why you got to a certain answer*
Answer #1

NaC3H5O2 (aq) -------------> Na^+ (aq)   + C3H5O2^- (aq)

0.022M                                                       0.022M

               C3H5O2^- (aq) + H2O(l) ------------------> C3H5O2H(aq)   +    OH^- (aq)

I               0.022                                                            0                          0

C              -x                                                                 +x                         +x

E             0.022-x                                                          +x                          +x

                   Kb     = Kw/Ka

                             = 1*10^-14/(1.34*10^-5)   = 7.46*10^-10

                       Kb     =   [C3H5O2H][OH^-]/[C3H5O2^-]

                      7.46*10^-10    = x*x/(0.022-x)

                     7.46*10^-10*(0.022-x)   = x^2

                           x   = 4.05*10^-6

              [OH^-]   =x    = 4.05*10^-6M

              POH   = -log[OH^-]

                        = -log(4.05*10^-6)

                          = 5.3925 >>>>>answer